What is the area of a rectangle with a perimeter of 18 units and a length of 5 units?

Answer by checkley79(3341)

P=2L+2W
18=2*5+2W
18=10+2W
2W=18-10
2W=8
W=8/2
W=4 ANS. FOR THE WIDTH.
AREA=LW
AREA=5*4=20 ANS. FOR THE AREA.